# itemizestditemize

### v1

@@ VIII Conclusion and Future Work
Similarly to $\kappa_{f}$, it also is $c_{f}\in[0,1]$. When $f$ is submodular, then $c_{f}=\kappa_{f}$. Generally, if $c_{f}=0$, then $f$ is modular, while if $c_{f}=1$, then eq. (34) implies the assumption that $f$ is non-decreasing. In [52], any monotone $f$ with total curvature $c_{f}$ is called $c_{f}$-submodular, as repeated below. [^7] Lehmann et al. [52] defined $c_{f}$-submodularity by considering in eq. (34) ${\cal A}\subseteq{\cal B}$ instead of ${\cal A}\subseteq{\cal V}$. Generally, non-submodular but monotone functions have been referred to as approximately or weakly submodular [53, 54], names that have also been adopted for the definition of $c_{f}$ in [52], e.g., in [55, 56].
Theorem 4 (Suboptimality Bounds for Submodular or Approximately Submodular Functions).
${\cal A}^{{\sf RAG}}$ enjoys the following approximation bounds:
  • if $f$ is non-decreasing submodular and ${\cal G}$ is fully centralized,
    $f(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,\frac{1}{2}\,f(\,\mathcal{A}^{\sf OPT}\,);$ (35)
  • if $f$ is non-decreasing submodular and ${\cal G}$ is not fully cent- ralized,
    $f(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,\frac{1-\kappa_{f}}{2-\kappa_{f}}\,f(\,\mathcal{A}^{\sf OPT}\,);$ (36)
  • if $f$ is non-decreasing and ${\cal G}$ is fully centralized,
    $f(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,\frac{1-c_{f}}{2-c_{f}}\,f(\,\mathcal{A}^{\sf OPT}\,);$ (37)
  • if $f$ is non-decreasing and ${\cal G}$ is not fully centralized,
    $f(\,{\cal A}^{{\sf RAG}}\,)\,\geq\,\frac{(1-c_{f})^{2}}{(1-c_{f})^{2}+1}\,f(\,\mathcal{A}^{\sf OPT}\,).$ (38)
@@
We present the proof separately for each case. First, when $f$ is submodular with ${\cal G}$ being fully centralized, RAG has the same $1/2$ bound as in Theorem 1 that follows from eqs. 28, 29 and 30.
@@
$\leq f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})$ (39)
$=\frac{2-\kappa_{f}}{1-\kappa_{f}}f({\cal A}^{{\sf RAG}})$ (40)
$\quad+\sum_{i\,\in\,{\cal N}}\,\left[f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})-\frac{1}{1-\kappa_{f}}f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]})\right]$
$\leq\frac{2-\kappa_{f}}{1-\kappa_{f}}f({\cal A}^{{\sf RAG}})$
$\quad+\sum_{i\,\in\,{\cal N}}\,\left[f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})-f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})\right]$ (41)
$=\frac{2-\kappa_{f}}{1-\kappa_{f}}f({\cal A}^{{\sf RAG}}),$ (42)
where eq. 39 follows eq. 22; eq. 40 holds by telescoping the sums; and eq. 41 holds from eq. 5:
$1-\kappa_{f}\leq\frac{f(\,a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}\setminus\{i\}}^{\sf RAG}\,)}{f(\,a_{i}^{\sf RAG}\,)}\leq\frac{f(\,a_{i}^{\sf RAG}\,|\,{\cal A}_{[i-1]}^{\sf RAG}\,)}{f(\,a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG}\,)}.$ (43)
When $f$ is non-submodular, and ${\cal G}$ is fully centralized,
@@
$=f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf OPT}\,|\,{\cal A}^{{\sf RAG}}\cup{\cal A}^{\sf OPT}_{[i-1]})$ (44)
$\leq f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,\frac{1}{1-c_{f}}f(a_{i}^{\sf OPT}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})$ (45)
$\leq f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,\frac{1}{1-c_{f}}f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})$ (46)
$=f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,\frac{1}{1-c_{f}}f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]})$ (47)
$=\frac{2-c_{f}}{1-c_{f}}f({\cal A}^{{\sf RAG}}),$ (48)
where eq. 44 follows from eq. 20, eq. 45 holds from Definition 4, eq. 46 holds since RAG selects $a^{{\sf RAG}}_{i}$ greedily, eq. 47 holds since ${\cal N}_{i}={\cal N}\setminus\{i\}$, and eq. 48 holds by telescoping the sums.
When $f$ is non-submodular, and ${\cal G}$ is not fully centralized,
$f({\cal A}^{\sf OPT})$
$\leq f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,\frac{1}{1-c_{f}}f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})$ (49)
$=\frac{(1-c_{f})^{2}+1}{(1-c_{f})^{2}}f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\frac{1}{1-c_{f}}f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})$
1($\hskip 99.58464pt-\frac{1}{(1-c_{f})^{2}}f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]})$ (50)
$\leq\frac{(1-c_{f})^{2}+1}{(1-c_{f})^{2}}f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\frac{1}{(1-c_{f})^{2}}f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]})$
1($\hskip 99.58464pt-\frac{1}{(1-c_{f})^{2}}f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]})$ (51)
$\leq\frac{(1-c_{f})^{2}+1}{(1-c_{f})^{2}}f({\cal A}^{{\sf RAG}}),$ (52)
where eq. 49 follows eq. 46, eq. 50 holds by telescoping the sums, and Proof of Theorem 4 holds from Definition 4. ∎
@@
$\delta_{i}^{\prime}({\cal S}\,|\,{\cal B}_{1})-\delta_{i}^{\prime}({\cal S}\,|\,{\cal B}_{1}\cup{\cal B}_{2})$ (55)
$=\delta_{i}^{\prime}({\cal S}\cup{\cal B}_{1})-\delta_{i}^{\prime}({\cal B}_{1})-\delta_{i}^{\prime}({\cal S}\cup{\cal B}_{1}\cup{\cal B}_{2})+\delta_{i}^{\prime}({\cal B}_{1}\cup{\cal B}_{2})$ (56)
@@
$\quad-f(\,a\,|\,{\cal A}_{{\cal S}\cup{\cal B}_{1}\cup{\cal B}_{2}}\,)+f(\,a\,|\,{\cal A}_{{\cal B}_{1}\cup{\cal B}_{2}}\,)$ (57)
$=f(\,a\cup{\cal A}_{{\cal S}\cup{\cal B}_{1}}\,)-f(\,{\cal A}_{{\cal S}\cup{\cal B}_{1}}\,)-f(\,a\cup{\cal A}_{{\cal B}_{1}}\,)+f(\,{\cal A}_{{\cal B}_{1}}\,)$
$\quad-f(\,a\cup{\cal A}_{{\cal S}\cup{\cal B}_{1}\cup{\cal B}_{2}}\,)+f(\,{\cal A}_{{\cal S}\cup{\cal B}_{1}\cup{\cal B}_{2}}\,)$
$\quad+f(\,a\cup{\cal A}_{{\cal B}_{1}\cup{\cal B}_{2}}\,)-f(\,{\cal A}_{{\cal B}_{1}\cup{\cal B}_{2}}\,)$ (58)
$=f(\,{\cal A}_{{\cal S}}\,|\,a\cup{\cal A}_{{\cal B}_{1}}\,)-f(\,{\cal A}_{{\cal S}}\,|\,{\cal A}_{{\cal B}_{1}}\,)$
$\quad-f(\,{\cal A}_{{\cal S}}\,|\,a\cup{\cal A}_{{\cal B}_{1}\cup{\cal B}_{2}}\,)+f(\,{\cal A}_{{\cal S}}\,|\,{\cal A}_{{\cal B}_{1}\cup{\cal B}_{2}}\,)$ (59)
$\leq f(\,{\cal A}_{{\cal S}}\,|\,a\cup{\cal A}_{{\cal B}_{1}}\,)-f(\,{\cal A}_{{\cal S}}\,|\,{\cal A}_{{\cal B}_{1}}\,)$
$\quad-f(\,{\cal A}_{{\cal S}}\,|\,a\cup{\cal A}_{{\cal B}_{1}}\,)+f(\,{\cal A}_{{\cal S}}\,|\,{\cal A}_{{\cal B}_{1}}\,)=0,$ (60)

### v2

@@ VIII Summary and Concluding Remarks
Similarly to $\kappa_{f}$, we have $c_{f}\in[0,1]$. When $f$ is submodular, then $c_{f}=\kappa_{f}$. Generally, if $c_{f}=0$, then $f$ is modular, while if $c_{f}=1$, then eq. (31) implies the assumption that $f$ is non-decreasing. In [59], any monotone $f$ with total curvature $c_{f}$ is called $c_{f}$-submodular, as repeated below. [^10] Lehmann et al. [59] defined $c_{f}$-submodularity by considering in eq. (31) ${\cal A}\subseteq{\cal B}$ instead of ${\cal A}\subseteq{\cal V}$. Generally, non-submodular but monotone functions have been referred to as approximately or weakly submodular [60, 61], names that have also been adopted for the definition of $c_{f}$ in [59], e.g., in [62, 63].
Theorem 4 (Suboptimality Bounds for Submodular and Approximately Submodular Functions).
If $f$ is not 2nd-order submodular, ${\cal A}^{{\sf RAG}}$ enjoys the following approximation bounds:
  • if $f$ is non-decreasing submodular, then
    $\frac{f(\,{\cal A}^{{\sf RAG}}\,)}{f(\,{\cal A}^{\sf OPT}\,)}\geq\begin{cases}\frac{1}{1+\kappa_{f}},&{\cal G}\text{ is {fully centralized}},\\ 1-\kappa_{f},&{\cal G}\text{ is {not fully centralized}},\end{cases}$ (32)
  • if $f$ is non-decreasing $c_{f}$-submodular, then
    $\frac{f(\,{\cal A}^{{\sf RAG}}\,)}{f(\,{\cal A}^{\sf OPT}\,)}\geq\begin{cases}\frac{1-c_{f}}{1+c_{f}-c_{f}^{2}},&{\cal G}\text{ is {fully centralized}},\\ (1-c_{f})^{2},&{\cal G}\text{ is {not fully centralized}}.\end{cases}$ (33)
@@
We present the proof separately for each case. First, when $f$ is submodular with ${\cal G}$ being fully centralized, RAG has the same $1/(1+\kappa_{f})$ bound as in Theorem 1, following from eq. 26.
@@
$f({\cal A}^{\sf OPT})$ $\leq f({\cal A}^{\sf RAG})+\kappa_{f}\sum_{i\in{\cal N}}f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})$
$\leq f({\cal A}^{\sf RAG})+\frac{\kappa_{f}}{1-\kappa_{f}}\sum_{i\in{\cal N}}f(a_{i}^{\sf RAG}\,|\,{\cal A}_{[i-1]}^{\sf RAG})$ (34)
$=\left(1+\frac{\kappa_{f}}{1-\kappa_{f}}\right)f({\cal A}^{{\sf RAG}}),$ (35)
where eq. 34 holds from eq. 25 and the definition of $\kappa_{f}$.
When $f$ is $c_{f}$-submodular, and ${\cal G}$ is fully centralized,
@@
$=f({\cal A}^{\sf OPT}\cup{\cal A}^{\sf RAG})-\sum_{i\in{\cal N}}f(a_{i}^{\sf RAG}\,|\,{\cal A}^{\sf OPT}\cup{\cal A}_{[i-1]}^{\sf RAG})$
$\leq f({\cal A}^{\sf RAG})+\frac{1}{1-c_{f}}\sum_{i\in{\cal N}}f(a_{i}^{\sf OPT}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})$
$\quad-(1-c_{f})\sum_{i\in{\cal N}}f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})$ (36)
$\leq f({\cal A}^{\sf RAG})+\left[\frac{1}{1-c_{f}}-(1-c_{f})\right]\sum_{i\in{\cal N}}f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})$ (37)
$=\left(c_{f}+\frac{1}{1-c_{f}}\right)f({\cal A}^{\sf RAG}),$ (38)
where eq. 36 holds from eq. 23 and Definition 5, eq. 37 holds since RAG selects $a^{{\sf RAG}}_{i}$ greedily, and eq. 38 holds since ${\cal N}_{i}\cap[i-1]=[i-1]$ for ${\cal G}$ being fully centralized.
When $f$ is $c_{f}$-submodular, and ${\cal G}$ is not fully centralized,
@@
$\leq f({\cal A}^{\sf RAG})+\left[\frac{1}{1-c_{f}}-(1-c_{f})\right]\sum_{i\in{\cal N}}f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})$
$\leq f({\cal A}^{\sf RAG})+\left[\frac{1}{(1-c_{f})^{2}}-1\right]\sum_{i\in{\cal N}}f(a_{i}^{\sf RAG}\,|\,{\cal A}_{[i-1]}^{\sf RAG})$ (39)
$=\frac{1}{(1-c_{f})^{2}}f({\cal A}^{\sf RAG}),$ (40)
where eq. 39 holds from eq. 37 and Definition 5. ∎
@@
$\delta_{i}^{\prime}({\cal S}\,|\,{\cal B}_{1})-\delta_{i}^{\prime}({\cal S}\,|\,{\cal B}_{1}\cup{\cal B}_{2})$
$=\delta_{i}^{\prime}({\cal S}\cup{\cal B}_{1})-\delta_{i}^{\prime}({\cal B}_{1})-\delta_{i}^{\prime}({\cal S}\cup{\cal B}_{1}\cup{\cal B}_{2})+\delta_{i}^{\prime}({\cal B}_{1}\cup{\cal B}_{2})$
@@
$\quad-f(a\,|\,{\cal A}_{{\cal S}\cup{\cal B}_{1}\cup{\cal B}_{2}})+f(a\,|\,{\cal A}_{{\cal B}_{1}\cup{\cal B}_{2}})\leq 0,$ (43)